OpenStudy (anonymous):

the government requires speedometers on cars sold in the united states to be accurate within +2.5% of the actual speed of the car. If your speedometer reads 60 miles per hour while you are driving on a highway , what is the range of possible actual speeds your car could be traveling at ?

OpenStudy (blues):

1. Calculate the range of 60 mph, + or - 2.5%. So the lower bound = 60 - 0.025*60 And the upper bound = 60 + 0.025*60